OpenStudy (anonymous):

There are 800 students in a school. If the boys were decreased by 80 and the girls increased by 25%, there would be an equal number of boys and girls. Find the original number of boys .

OpenStudy (anonymous):

let the number of girls be x and that of boys be 800-x by th eproblem 800-x-80 = (100+25)% x 720-x=1.25x or 2.25x=720 hence x=720/2.25 or x=320 hence number of boys =800-320=480